X=5t^2 - 2t. Find the velocity at t=3s

tX=5


1
Expert's answer
2021-04-29T07:27:48-0400

"x(t) = 5t^2-2t;"

"V(t) = x'(t)= 10t-2;"

"V(3)= 10*3-2 = 28"

"V(5) = 10*5 -2 = 48"

"Answer: V(3) = 28;V(5) =48"